OpenStudy (swissgirl):

Prove that if f: A -->B and g: B-->C and g(f(x)): A -->C which is one to one then f: A --> B is one to one

OpenStudy (anonymous):

work directly from the definition of one to one

OpenStudy (swissgirl):

if f(x)=f(y) then x=y

OpenStudy (anonymous):

that is, start with \(g(f(a))=g(f(b))\implies a=b\)

OpenStudy (anonymous):

now suppose \(f\) is not one one one then there exists \(a\neq a'\in A\) such that \(f(a)=f(a')\)

OpenStudy (swissgirl):

Riighhhttt

OpenStudy (anonymous):

but if this is true, then since \(f(a)=f(a')\) it must be true that \(g(f(a))=g(f(a'))\) since \(g\) is a function

OpenStudy (anonymous):

contradicting the assumption

OpenStudy (swissgirl):

alrighty gotchhhaaaaaaaa

OpenStudy (swissgirl):

Thanks :)